OpenStudy (anonymous):

integral from x to lnx (t^2 + 3)dt Please show steps.

OpenStudy (kirbykirby):

Ok give me a sec to type it out

OpenStudy (anonymous):

Awesome, thanks.

OpenStudy (anonymous):

did you mean "t" to "ln(t)" ? otherwise T is just a constant

OpenStudy (anonymous):

Yes

OpenStudy (kirbykirby):

\[\int\limits_{X}^{\ln X}(t^2+3)dt\]=\[[\frac{t^3}{3}+3t] \]evaluated from x to ln x (this is just the power rule) = \[(\frac{(\ln x)^3}{3}+3(\ln x))-(\frac{x^3}{3}+3x)\] = \[\frac{(\ln x)^3}{3}+3(\ln x)-\frac{x^3}{3}-3x\]

OpenStudy (anonymous):

Pretty good @kirbykirby

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (kirbykirby):

yw